Two sides of the triangle are 26 and 27, while the height to the third side is 25. Find the area of the triangle. Please explain.
Answer by rothauserc(4718)

the height forms two right triangles, let their bases be x and y
:
25^2 +x^2 = 26^2
:
x^2 = 26^2 - 25^2 = 51
:
x = 7.14 approximately 7.1
:
25^2 +y^2 = 27^2
:
y^2 = 27^2 - 25^2 = 104
:
y = 10.19 approximately 10.2
:
Base of triangle = 7.1 + 10.2 = 17.3
:
Area of triangle = (1/2) * 17.3 * 25 = 216.25 square units
: